OpenStudy (anonymous):

The first four terms of a sequence are shown below: 8, 5, 2, -1. Which of the following functions best defines this sequence?

OpenStudy (kiakat):

what are the functions??

OpenStudy (anonymous):

see the first term is 8 and 2nd term-1st term=3rd term -2nd term=...=-3 hence it is an AP with common difference -3 the series is given by a+(n-1)d=8+(n-1)(-3)=-3n+11 for all n >=1

OpenStudy (anonymous):

@kgdl